Grade 11General Physics


The ratio of the times taken by a body moving with uniform acceleration in reaching two points P and Q along a straight line path 1:2. If the body starts from rest and moves linearly, the ratio of the distances of P and Q from the starting point is

Saksham Bhatnagar
10 Years agoGrade 11

1 Answer

Beenu Mathew
10 Years ago
starting from rest (u=0)
v=at (where v is final veocity, a is acceleration and t is time)
you could do this..
t1/t2 = ½ = a2 /a1 
ie a1/a2=2
now 
s= ½ at2
s1/s2 = ½ ( on solving)
HOPE THIS HELPS
BEST OF LUCK!!